# One-Versus-Others Attention: Scalable Multimodal Integration for Clinical Data

### v3

@@ 4.1.3 MIMIC-IV and CXR
MIMIC-IV (Johnson et al., 2023) covers 431K visits for 180K patients admitted to the ICU in the Beth Israel Deaconess Medical Center. MIMIC Chest X-ray (MIMIC-CXR) (Johnson et al., 2019) contains chest radiographs in DICOM format with free-text radiology reports. The dataset contains 377,110 images corresponding to 227,835 radiographic studies performed at Beth Israel Medical Center. We follow the pre-processing of MedFuse (Mohammed et al., 2021) to extract the clinical time-series data from MIMIC-IV along with the associated chest X-ray images in MIMIC-CXR. We further expand the number of modalities by adding a demographics table and discharge notes, resulting in four modalities. We also follow MedFuse in the construction of the phenotyping task. The goal of this multi-label classification task is to predict whether a set of 25 chronic, mixed, and acute care conditions are assigned to a patient in a given ICU stay. This is a 25-class multi-label task with four modalities.
4.1.4 eICU
@@
The eICU collaborative database encompasses data derived from intensive care units (ICU) across the continental United States between 2014 and 2015 (Pollard et al., 2018). The database is structured with a collection of tables relating to patients in the program, linked through a patient unit stay ID. For our predictive task aimed at determining mortality, we focus on six tabular modalities represented by the patient, diagnosis, treatment, medication, lab, and apacheApsVar tables. The apacheApsVar contains numerical variables used to calculate the Acute Physiology Score (APS) for patients, which summarizes a patient’s severity of illness. Details on other modalities can be found in Appendix C. After isolating ICU stays that had complete data for all modalities, our dataset comprises 75,845 unique patients associated with 93,784 ICU stays, of which 86,012 were recorded as alive and 7,772 were recorded as dead. This is a 2-class classification task with six modalities.
@@ 4.2 Baselines
Our multimodal baselines include a conventional concatenation fusion with no attention, early fusion followed by self-attention, and pairwise cross-attention fusion. The architectures of all models are identical except for their integration stage. For example, since modality-specific encoders can produce different dimension sizes, we add a linear layer before integration to create the same input dimensions. While this step is not strictly necessary for concatenation, we still add the layer there so that no additional factors influence computation costs and performance. While there are many multimodal Transformers available for the vision-language domain, our focus is on examining the underlying fusion mechanism and creating a general integration paradigm for any application, especially ones outside of vision-language.
@@ 4.3 Implementation details
For the MIMIC dataset, we follow the established train, validation, and test split in (Hayat et al., 2022). Similarly, for the TADPOLE task, we use the provided data splits, but add a constraint that repeating patients cannot appear across data splits, as to avoid information leakage. In the other datasets, for consistency, we randomly sampled 80% of the data for the training set and 10% each for test and validation sets, as there was not an established and publicly accepted split. To evaluate our model against other integration techniques, we use the domain-accepted metrics for each task: For MIMIC and eICU, we use area under the receiver operating characteristic (AUROC) and area under Precision-Recall (PR) curve (AUPRC) as established in past works (Hayat et al., 2022; Sheikhalishahi et al., 2020); For TADPOLE we use the multi-class area under the receiver operating curve (mAUC) and the overall balanced classification accuracy (BCA), as established by the competition creators (Marinescu et al., 2019); For TCGA, we use accuracy and F1-score as measures of performance, used in (Mohammed et al., 2021). For all datasets, we used the number of floating-point operations (FLOPs) as the measure of runtime complexity. FLOPs were measured per sample and reported as the difference between concatenation, the simplest integration setting, and multimodal attention (FLOPs).
@@ 4.4 Hyperparameter Tuning
Our hyperparameter tuning scheme was consistent for each dataset and each model. For each experiment, we used the evaluation metrics on the validation set to determine the best hyperparameters. We tuned the learning rate (0.01 - 1 x , dividing by 10 for each interval), batch size (16, 32, 64, 128), epochs (200 epochs with early stopping if validation performance did not increase for 5 epochs), and number of attention heads for OvO, self-attention, and cross-attention models (1, 2, 4, 8, 16). For the neural network encoders, we tuned the number of linear layers ranging from 1 to 4. Similarly, for the convolutional neural network, we tuned the number of convolution layers ranging from 1 to 4. For compute times and GPU details used for hyperparameter tuning, see Appendix G. Lastly, we randomly picked 10 random seeds for every experiment - once the best hyperparameters were picked, ten models initialized with those seeds and parameters were run. Then, using the trained ten models, we evaluated on the test set and took the average of the 10 runs along with the standard deviation, which is reported in Section 5.
@@ 5 Results
Using four real-world clinical datasets, diverse in terms of the number of modalities, feature space, and classification tasks, we demonstrate that our method consistently and drastically reduces computational costs compared to early fusion and pairwise fusion while simultaneously maintaining or enhancing performance. This is demonstrated on four-modality, five-modality, and two six-modality datasets.
For the four-modality MIMIC task, we used a pre-trained ClinicalBERT model for the text modality. For all other modalities, we used the appropriate neural network architecture (i.e., CNN for images, LSTM for time series, and a multi-layer perception for all tabular data). We perform significance testing between OvO attention’s and the next best performing model, detailed in the Appendix H.
The results on MIMIC are presented in Table 2, clearly demonstrating the scalability and performance advantages of OvO attention. OvO’s 4,227,072 FLOPs notably reduce computational costs compared to cross-attention (52,723,712 FLOPs) and self-attention (67,633,152 FLOPs), achieving reductions by 91.98% and 93.75%, respectively, thus highlighting OvO’s superior efficiency. The unimodal results show that the textual modality is most valuable in phenotype prediction, and ClinicalBERT alone performs better than self-attention and cross-attention. This indicates that the added complexity and forced interactions are not necessarily conducive to result quality. However, OvO attention can extract information from the other modalities for a significant performance increase rather than a decrease (p-value <0.01, see Appendix H).
| Model | Modalities | FLOPs | AUROC | AUPRC |
| LSTM | TS | - | 58.8 ±0.6 | 28.5 ±0.4 |
| CNN | I | - | 56.9 ±0.3 | 26.7 ±0.2 |
| NN | D | - | 64.1 ±0.4 | 32.4 ±0.3 |
| ClinicalBERT | T | - | 79.3 ±0.4 | 58.7 ±0.3 |
| Concatenation | All | * | 82.7 ±0.6 | 65.1 ±1.8 |
| Cross-Attention | All | 52,723,712 | 78.2 ±2.1 | 54.1 ±2.7 |
| Self-Attention | All | 67,633,152 | 78.5 ±2.0 | 55.7 ±3.1 |
| OvO Attention | All | 4,227,072 | 83.6 ±1.1 | 66.2 ±2.6 |
For the five-modality cancer prediction task from TCGA, we show our results in Table 3. While here, cross-attention yielded slightly better performance than OvO, our model offers substantial scalability benefits and, across 10 random seeds, performed consistently, showing robustness and stability. Furthermore, the difference between cross-attention and OvO attention accuracy and F1-scores was not statistically significant (p-value >0.01, see Appendix H). More importantly, OvO’s 330,240 FLOPs significantly undercut cross (5,260,800 FLOPs) and self-attention (6,556,160 FLOPs), achieving reductions of 93.73% and 94.96%, respectively, highlighting OvO’s superior efficiency.
| Model | Modalities | FLOPs | Accuracy | F1-Score |
| CNN | I | - | 56.7 ±3.0 | 55.6 ±4.0 |
| NN | C | - | 60.8 ±0.5 | 57.2 ±0.3 |
| NN | CNV | - | 93.4 ±0.8 | 94.1 ±0.7 |
| NN | M | - | 97.0 ±0.4 | 97.4 ±0.4 |
| NN | GE | - | 97.6 ±1.2 | 97.7 ±1.0 |
| Concatenation | All | * | 97.8 ±1.5 | 97.9 ±1.3 |
| Self-Attention | All | 6,556,160 | 96.8 ±1.0 | 97.0 ±0.9 |
| Cross-Attention | All | 5,260,800 | 99.2 ±1.0 | 99.3 ±0.9 |
| OvO Attention | All | 330,240 | 98.3 ±0.1 | 98.4 ±0.1 |
@@
| Model | Modalities | FLOPs | MAUC | BCA |
| NN | A | - | 63.5 ±3.1 | 56.4 ±3.8 |
| NN | CSF | - | 64.4 ±1.1 | 53.6 ±2.7 |
| NN | M | - | 67.0 ±1.3 | 57.2 ±1.0 |
| NN | F | - | 66.6 ±0.3 | 60.8 ±0.7 |
| NN | D | - | 74.6 ±0.9 | 62.0 ±0.6 |
| NN | CT | - | 97.8 ±0.2 | 88.6 ±0.7 |
| Concatenation | All | * | 97.7 ±0.8 | 91.9 ±1.9 |
| Cross-Attention | All | 8,921,088 | 97.1 ±0.6 | 90.7 ±1.7 |
| Self-Attention | All | 9,633,792 | 94.8 ±1.1 | 86.6 ±2.6 |
| OvO Attention | All | 405,504 | 98.3 ±0.4 | 93.0 ±1.4 |
@@
Lastly, the results on the six-modality eICU mortality prediction task are shown in Table 5, demonstrating the scalability and performance advantages of OvO attention. OvO’s 6,340,608 FLOPs significantly undercut those of cross-attention (129,957,888 FLOPs) and self-attention (151,781,376 FLOPs), achieving reductions of approximately 95.12% and 95.82%, respectively, thereby highlighting OvO’s efficiency. Mirroring the trends observed in the MIMIC and TADPOLE datasets, we note a dominant unimodal modality, specifically Lab modality, in this experiment as well. While concatenating modalities does enhance performance, this improvement is not seen in self and cross-attention models. In contrast, OvO attention not only reflects these performance gains but does so significantly (p-value <0.01). We hypothesize that this is due to the overfitting of more complex integration frameworks of self and cross-attention on relatively smaller clinical datasets. OvO, in its simplicity akin to concatenation, manages to strike a balance by maintaining flexibility and capturing inter-modal interactions through its attention mechanism, thus offering an edge in performance without excessive complexity.
| Model | Modalities | FLOPs | AUROC | AUPRC |
| NN | DE | - | 50.2 ±0.6 | 91.8 ±0.2 |
| NN | M | - | 56.3 ±1.3 | 93.1 ±0.3 |
| NN | D | - | 58.2 ±2.1 | 93.3 ±0.4 |
| NN | T | - | 66.1 ±0.5 | 94.8 ±0.1 |
| NN | A | - | 77.6 ±0.2 | 97.0 ±0.1 |
| NN | L | - | 81.5 ±0.4 | 97.0 ±0.1 |
| Concatenation | All | * | 81.7 ±1.6 | 97.5 ±0.3 |
| Cross-Attention | All | 129,957,888 | 77.6 ±1.6 | 95.4 ±0.3 |
| Self-Attention | All | 151,781,376 | 80.2 ±2.0 | 96.8 ±0.4 |
| OvO Attention | All | 6,340,608 | 82.5 ±0.9 | 97.8 ±0.2 |
@@
In summary, across diverse clinical datasets and modalities, OvO attention consistently outperforms traditional fusion techniques in both predictive performance and computational efficiency, underlining its robustness and adaptability in handling complex multimodal healthcare data.
5.1 Generalizability
While our work focused on the clinical domain, where OvO will be most impactful and relevant, our method is a domain-agnostic approach and can be highly effective in other multimodal scenarios. We use the Amazon reviews dataset (Ni et al., 2019), to demonstrate OvO’s performance on a non-clinical dataset. The Amazon Reviews dataset includes review images, review text, and product metadata (e.g., price, category), with the goal of review sentiment classification (positive or negative). See Appendix F for more pre-processing details. The Amazon Reviews results are shown in Table 6 and demonstrate both the scalability and performance advantages of OvO attention. OvO’s 523,520 FLOPs significantly undercut those of cross-attention (1,903,616 FLOPs) and self-attention (2,685,440 FLOPs), achieving reductions of 72.50% and 80.51%, respectively, thereby highlighting OvO’s efficiency on the integration. Even though the efficiency gains scale with the number of modalities (see Figure 2), it is impressive to see significant improvements can be achieved for a dataset with just three modalities. Since the textual modality is most valuable in sentiment prediction, the performance of BERT alone is higher than concatenation, self-attention, and cross-attention. This indicates that the noise from metadata and images interferes with model performance. However, OvO attention can extract information from the other two modalities for a significant performance increase rather than a decrease (p-value <0.01).
| Model | Modalities | FLOPs | Accuracy | F1-Score |
| NN | Tb | - | 57.6 ±0.7 | 57.5 ±0.8 |
| ResNet | I | - | 66.3 ±0.7 | 66.6 ±0.6 |
| BERT | T | - | 92.6 ±0.5 | 92.9 ±0.4 |
| Concatenation | All | * | 92.2 ±0.4 | 92.8 ±0.3 |
| Self-Attention | All | 2,685,440 | 92.4 ±0.4 | 92.4 ±0.4 |
| Cross-Attention | All | 1,903,616 | 91.6 ±0.7 | 92.2 ±0.6 |
| OvO Attention | All | 523,520 | 93.1 ±0.3 | 93.0 ±0.3 |
Overall, the adaptability and performance enhancement demonstrated by OvO across various applications, not limited to the medical field, highlight its potential as a universal scalable solution in multimodal learning. OvO’s domain-agnostic approach can significantly contribute to overcoming the computational bottlenecks commonly encountered in multimodal scenarios.
6 Discussion
Our proposed method, OvO, provides a way to overcome one of the major challenges associated with multimodal datasets - computational resource demand and cost. Thus enabling adoption in resource-constrained domains, such as clinical decision support.
The literature on multimodal models and, specifically, Multimodal Transformers provides different architectures, some of which make efforts towards scalability. However, such models focus on the vision-language domain, and are not applicable in all multimodal scenarios (e.g., clinical decision support). Thus, we chose common fusion paradigms (similar to (Nagrani et al., 2021)) that encapsulate most of the underlying attention mechanisms in the multimodal literature. We aim to modify the basic attention mechanism in such models to make multimodal fusion more scalable with an increasing number of modalities.
@@ 7 Conclusion
We present One-Versus-Others (OvO), a new scalable multimodal attention mechanism. The proposed formulation significantly reduces the computational complexity compared to the widely used early fusion through self-attention and cross-attention methods. Notably, OvO achieves, at minimum, a reduction of 91.98% in FLOPs when benchmarked against self and cross-attention methods across a range of clinical datasets encompassing four, five, and six modalities. We provide both a detailed theoretical complexity analysis and empirical evidence from a simulated experiment, illustrating that OvO’s computational demand scales linearly with the number of modalities, in contrast to the quadratic scaling observed in other methods. Furthermore, the versatility of OvO is evidenced through its application to a non-clinical dataset, where it consistently surpassed all comparative baselines. Overall, the results unequivocally establish that OvO not only significantly reduces computational expenses but also meets or exceeds the performance of existing state-of-the-art fusion methodologies.
8 Impact Statement
Our proposed work, OvO, provides a method to overcome one of the major challenges associated with multimodal datasets - computational resource demand and cost. The scalability of the proposed formula allows diverse multimodal experiments to run in a wider range of research environments despite economic and computational constraints. While our method of modality fusion is meant to reduce the negative environmental impact of attention-learning models, we note that the individual modality encoders used (e.g., ClinicalBERT) still have a negative impact by creating emissions.
@@ A.1 Early Fusion
Attention scores are computed by taking the dot product of queries and keys. The self-attention mechanism has quadratic complexity with respect to the sequence length and linear complexity with respect to the representation dimension [Vaswani et al., 2017]. Thus, given the concatenated sequence’s length of and the dimension of the keys and queries , the complexity of this step is:

### v4

@@ 4.1.1 MIMIC-IV and CXR data
MIMIC-IV [9] covers 431K visits for 180K patients admitted to the ICU in the Beth Israel Deaconess Medical Center. MIMIC Chest X-ray (MIMIC-CXR) [8] contains chest radiographs in DICOM format with free-text radiology reports. The dataset contains 377,110 images corresponding to 227,835 radiographic studies performed at Beth Israel Medical Center. We follow the pre-processing of MedFuse [19] to extract the clinical time-series data from MIMIC-IV along with the associated chest X-ray images in MIMIC-CXR. We further expand the number of modalities by adding a demographics table and discharge notes, resulting in four modalities. We also follow MedFuse in the construction of the phenotyping task. The goal of this multi-label classification task is to predict whether a set of 25 chronic, mixed, and acute care conditions are assigned to a patient in a given ICU stay. This is a 25-class multi-label task with four modalities.
4.1.2 The Alzheimer’s Disease Prediction Of Longitudinal Evolution (TADPOLE) data
The Alzheimer’s Disease Neuroimaging Initiative (ADNI) [21] database provides neuroimaging data, cognitive test scores, biomarker profiles, and genetic information for Alzheimer’s disease (AD), mild cognitive impairment (MCI), and normal patients. We use the processed data from the Alzheimer’s Disease Prediction Of Longitudinal Evolution (TADPOLE) challenge [17]. We focus on a one-time diagnosis prediction task, using the most recent available data point for each patient across all modalities. This ensures that each data sample includes information from the same time point, which aligns with our goal of evaluating cross-modal integration rather than longitudinal progression. We utilize six modalities that have the least missing information per patient: cognitive tests - neuropsychological tests administered by a clinical expert; MRI ROIs (generated from Freesurfer) - measures of brain structural integrity; FDG PET ROI averages - measure cell metabolism, where cells affected by AD show reduced metabolism; AV45 PET ROI averages - measures amyloid-beta load in the brain; demographic information (e.g., age, gender, education); and CSF biomarkers - amyloid and tau levels in the cerebrospinal fluid. The preprocessing provided by TADPOLE turned every modality into a tabular form (including imaging). After removing patients with missing modalities, we had 767 MCI patients, 493 normal patients, and 143 AD patients. Thus, this is a three-class classification task with six modalities.
4.1.3 eICU data
@@ 4.1.2 The Alzheimer’s Disease Prediction Of Longitudinal Evolution (TADPOLE) data
The eICU collaborative database includes data from ICUs across the continental United States between 2014 and 2015 [22]. It consists of tables linked through a patient unit stay ID. For our mortality prediction task, we focus on six tabular modalities: patient, diagnosis, treatment, medication, lab, and apacheApsVar tables. The apacheApsVar table contains numerical variables used to calculate the Acute Physiology Score (APS), an established method within the Acute Physiology Age Chronic Health Evaluation (APACHE) system for summarizing patient’s severity of illness on ICU admission and predicting outcomes. The patient table includes demographic, admission, and discharge details, and is used to determine mortality status. The diagnosis table lists active diagnoses for each patient, the treatment table includes active treatments, and the medication table contains active medication orders. We extract the features from these tables by one-hot encoding the relevant conditions, treatment types, and drug names, respectively. The lab table includes lab results, with features extracted by summing commonly recorded lab types. Our dataset includes 75,845 unique patients with 93,784 ICU stays, 86,012 recorded as alive and 7,772 as dead. This is a 2-class classification task with six modalities.
@@ 4.2 Baselines
Our multimodal baselines include a conventional concatenation fusion with no attention, early fusion followed by self-attention, and pairwise cross-attention fusion. The architectures of all models are identical except for their integration stage. For example, since modality-specific encoders can produce different dimension sizes, we add a linear layer before integration to create the same input dimensions. Although this step is not strictly necessary for concatenation, we still add the layer there so that no additional factors influence computation costs and performance. While there are many multimodal Transformers available for the vision-language domain, our focus is on examining the underlying fusion mechanism and creating a general integration paradigm for any application, especially ones outside of vision-language. In Appendix Section 5, we touch on the limitations of our experiments and future work that we did not cover.
@@ 4.3 Implementation details
For the MIMIC dataset, we follow the established train, validation, and test split in Hayat et al [6]. Similarly, for the TADPOLE task, we use the provided data splits but add a constraint that repeating patients cannot appear across data splits to avoid information leakage. In the other datasets, for consistency, we randomly sampled 80% of the data for the training set and 10% each for test and validation sets, as there was not an established split. To evaluate our model against other integration techniques, we use the domain-accepted metrics for each task: For MIMIC and eICU, we use area under the receiver operating characteristic (AUROC) and area under Precision-Recall (PR) curve (AUPRC) as established in past works [6, 28]; For TADPOLE we use the multi-class area under the receiver operating curve (mAUC) and the overall balanced classification accuracy (BCA), as established by the competition creators [17]. For all datasets, we used the number of floating-point operations (FLOPs) as the measure of runtime complexity. FLOPs were measured per sample and reported as the difference between concatenation, the simplest integration setting, and multimodal attention (FLOPs).
@@ 4.4 Hyperparameter Tuning
Our hyperparameter tuning scheme was consistent for each dataset and each model. For each experiment, we used the evaluation metrics on the validation set to determine the best hyperparameters. We tuned the learning rate (0.01 - 1 x , dividing by 10 for each interval), batch size (16, 32, 64, 128), epochs (200 epochs with early stopping if validation performance did not increase for 5 epochs), and number of attention heads for OvO, self-attention, and cross-attention models (1, 2, 4, 8, 16). For the neural network encoders, we tuned the number of linear layers ranging from 1 to 4. Similarly, for the convolutional neural network, we tuned the number of convolution layers ranging from 1 to 4. For compute times and GPU details used for hyperparameter tuning, see Appendix Section 2. Lastly, we randomly picked 10 random seeds for every experiment - once the best hyperparameters were picked, ten models initialized with those seeds and parameters were run. Then, using the trained ten models, we evaluated on the test set and took the average of the 10 runs along with the standard deviation, which is reported in Section 5.
@@ 5 Results
Using three real-world clinical datasets, diverse in terms of the number of modalities, feature space, and classification tasks, we demonstrate that our method consistently and drastically reduces computational costs compared to early fusion and pairwise fusion while simultaneously maintaining or enhancing performance. This is demonstrated on one four-modality dataset and two six-modality datasets.
For the four-modality MIMIC task, we used pre-trained ClinicalBERT model for the text modality and fine-tuned it for the unimodal baseline and the multimodal task, separately, ensuring adaptation in each setting. For all other modalities, we used the appropriate neural network architecture (i.e., CNN for images, LSTM for time series, and a multi-layer perception for all tabular data). We perform significance testing between OvO attention and the next best-performing model, detailed in Appendix Section 3.
[…]
The results on MIMIC are presented in Table 5, clearly demonstrating the scalability and performance advantages of OvO attention. OvO’s 4,227,072 FLOPs notably reduce computational costs compared to cross-attention (52,723,712 FLOPs) and self-attention (67,633,152 FLOPs), achieving reductions by 91.98% and 93.75%, respectively, thus highlighting OvO’s superior efficiency. The unimodal results show that the textual modality is most valuable in phenotype prediction, and ClinicalBERT alone performs better than self-attention and cross-attention. This indicates that the added complexity and forced interactions are not necessarily conducive to result quality. However, OvO attention can extract information from the other modalities for a significant performance increase rather than a decrease (p-value <0.01, see Appendix Section 3).
@@
Lastly, the results on the six-modality eICU mortality prediction task are shown in Table 5, demonstrating the scalability and performance advantages of OvO attention.
[…]
@@
OvO’s 6,340,608 FLOPs significantly undercut those of cross-attention (129,957,888 FLOPs) and self-attention (151,781,376 FLOPs), achieving reductions of approximately 95.12% and 95.82%, respectively, thereby highlighting OvO’s efficiency. Mirroring the trends observed in the MIMIC and TADPOLE datasets, we note a dominant unimodal modality, specifically Lab modality, in this experiment as well. While concatenating modalities does enhance performance, this improvement is not seen in self and cross-attention models. In contrast, OvO attention not only reflects these performance gains but does so significantly (p-value <0.01). We hypothesize that this is due to the overfitting of more complex integration frameworks of self and cross-attention on relatively smaller clinical datasets. OvO, in its simplicity akin to concatenation, manages to strike a balance by maintaining flexibility and capturing inter-modal interactions through its attention mechanism, thus offering an edge in performance without excessive complexity.
@@
In summary, across diverse clinical datasets and modalities, OvO attention consistently outperforms traditional fusion techniques in both predictive performance and computational efficiency, underlining its robustness in handling complex multimodal healthcare data.
@@ 6 Conclusion
We present One-Versus-Others (OvO), a new scalable multimodal attention mechanism. The proposed formulation significantly reduces the computational complexity compared to the widely used early fusion through self-attention and cross-attention methods. Notably, OvO achieves, at minimum, a reduction of 91.98% in FLOPs when benchmarked against self and cross-attention methods across a range of clinical datasets containing up to six modalities. We provide both a detailed theoretical complexity analysis and empirical evidence from a simulated experiment, illustrating that OvO’s computational demand scales linearly with the number of modalities, in contrast to the quadratic scaling observed in other methods. Our proposed method provides a way to overcome one of the major challenges associated with multimodal datasets - computational resource demand and cost, thus enabling adoption in resource-constrained domains, such as clinical decision support. Overall, the results unequivocally establish that OvO not only significantly reduces computational expenses but also exceeds the performance of existing state-of-the-art fusion methodologies.
@@ A.1 Early Fusion
Attention scores are computed by taking the dot product of queries and keys. The self-attention mechanism has quadratic complexity with respect to the sequence length and linear complexity with respect to the representation dimension [32]. Thus, given the concatenated sequence’s length of and the dimension of the keys and queries , the complexity of this step is: